# Spatial segregation of massive clusters in a simulation of colliding dwarf galaxies

### v1

@@ Abstract
The collective properties of star clusters are investigated using a simulation of the collision between two dwarf galaxies. The characteristic power law of the cluster mass function, $N(M)$, with a slope $d\log N/d\log M\sim-1$, is present from cluster birth and remains throughout the simulation. The maximum mass of a young cluster scales with the star formation rate (SFR). The relative average minimum separation, ${\cal R}(M)=N(M)^{1/p}{\bar{D}}_{\rm min}(M)/{\bar{D}}(M_{\rm low})$, for average minimum distance ${\bar{D}}_{\rm min}(M)$ between clusters of mass $M$, and for lowest mass, $M_{\rm low}$, measured in projection ($p=2$) or three dimensions ($p=3$), has a negative slope, $d\log{\cal R}/d\log M\sim-0.2$, for all masses and ages. This agrees with observations of ${\cal R}(M)$ in low-mass galaxies studied previously. Like the slope of $N(M)$, ${\cal R}(M)$ is apparently a property of cluster birth for dwarf galaxies that does not depend on SFR or time. The negative slope for ${\cal R}(M)$ implies that more massive clusters are centrally concentrated relative to lower mass clusters throughout the entire mass range. Cluster growth through coalescence is also investigated. The ratio of the kinetic to potential energy of all near-neighbor clusters is generally large, but a tail of low values in the distribution of this ratio suggests that a fraction of the clusters merge, $\sim 8$% by number throughout the $\sim 300$ Myr of the simulation and up to 60% by mass for young clusters in their first 10 Myr, scaling with the SFR above a certain threshold.
@@ 1 Introduction
Stars form in dense interstellar gas that has a hierarchical structure from turbulence and self-gravity (e.g., Scalo, 1985; Stanimirovic et al., 1999; Elmegreen et al., 2001). Young stellar positions are similarly hierarchical with star clusters forming in the mixed inner parts of the densest regions (e.g., Zhang et al., 2001; Elmegreen, 2010; Guszejnov et al., 2022). For a self-similar hierarchy, the mass distribution function of concentrations has a power law with a slope of around $-2$ (Fleck, 1996; Elmegreen & Falgarone, 1996), which explains the power law portion of the mass function that is commonly observed for clusters (Elmegreen & Efremov, 1997; Zhang et al., 1999; Krumholz et al., 2019; Cook et al., 2019). The power-law portion of the stellar mass function may have a similar origin (Elmegreen, 1997; Hennebelle & Chabrier, 2008; Shadmehri & Elmegreen, 2011).
Power laws from structure rather than self-regulated processes have the property that the maximum mass of an object that forms scales with the total number of objects (Elmegreen, 1983). Such scaling has been observed for both star clusters (Larsen, 2002; Billett et al., 2002; Whitmore, 2002; Hunter et al., 2003; Bastian, 2008; Cook et al., 2023) and stars (Corbelli et al., 2009; Andrews et al., 2014; Jung et al., 2023).
@@ 3 Cluster Mass Functions
Figure 2: The mass functions of clusters at the 5 fiducial times in increasing order of the colors: blue, red, green, black, cyan. The mass functions for all clusters are on the left and for clusters younger than 10 Myr are on the right. Power-law fits to the declining parts of the functions are shown in the same colors. The slopes are all around $-1$, as indicated in Table 1.
@@
Figure 2 plots the cluster mass functions at the five fiducial times in the following sequence of colors: blue for the first time, 78 Myr, followed by red (128 Myr), green (169 Myr), black (211 Myr), and cyan for the last time, 277 Myr. The mass functions for all the clusters regardless of age are on the left and for clusters younger than 10 Myr are on the right. The straight lines are linear fits (on these log-log plots) to the right of the peaks. The slopes $\alpha$ of these lines are the power law indices of the mass functions. They are tabulated in Table 1. For a mass function written in linear intervals of mass, the negative slope is 1 greater than $\alpha$, i.e., $n(M)\propto M^{-\alpha+1}$ for masses between $M$ and $M+dM$. The highest-mass cluster is the single peak on the far right of these plots; it appears around the time of the peak in the SFR.
@@
Time (Myr) Slope intercept
@@
Figure 3: The slopes $\alpha$ of the mass functions for all clusters (blue) and young clusters (red) are shown versus time. They hover around $\alpha=-0.8$ to $-1$.
@@
Figure 6: The slope of the power-law relation between the relative average minimum separation and the cluster mass, $d\log{\cal R}(M)/d\log M$, is shown versus time for all clusters (top) and young clusters (bottom) and for projected, 2D (red) and 3D (blue) separations. The slope hovers around $-0.2$.
For a mass function slope of $\alpha=-1$, the largest cluster mass is proportional to the number of clusters. Figure 4 shows the maximum cluster mass versus time in the left-hand panel, total stellar and cluster mass in the middle panel, and SFR in the right-hand panel. In the left and right-hand panels, the blue curves are for all cluster ages and the red curves are for clusters younger than 10 Myr. The maximum mass of a young cluster follows the SFR even as the SFR increases and decreases. The maximum mass for all ages does not follow the varying SFR because this mass does not decrease when the SFR drops.
@@ 4 Relative Average Minimum Separation and Projected Separation
Figure 7: The slope of the power-law relation between the relative average minimum separation and the cluster mass, $d\log{\cal R}(M)/d\log M$, is shown versus cluster age for the time interval between 250 Myr and 299 Myr. Two-dimensional and three-dimensional cluster separations are shown in red and blue.
@@
Figure 5 shows ${\cal R}(M)$ for 3D (top) and 2D (bottom) distributions, and for all clusters (left) and clusters younger than 10 Myr (right). In each panel, ${\cal R}(M)$ is given at the 5 fiducial times with the same color code as in Figure 2. ${\cal R}(M)$ decreases with mass once there are enough clusters to obtain a meaningful slope, which is starting with the main starburst and for times of 169 Myr or later in the figure. Because both 3D and 2D curves have their minimum average separations ${\bar{D}}$ normalized at the lowest masses, their absolute values in these plots are proportional to $N(M)$ to the corresponding powers $1/3$ or $1/2$, making the 2D curves higher.
@@
Figure 6 shows the slope $d\log{\cal R}(M)/d\log M$ versus time for 3D (blue) and 2D (red) distributions and for all clusters (top) and young clusters (bottom). Once there are enough clusters to have a reasonable slope, this slope is clearly negative even for young clusters. The 2D slope is slightly steeper than the 3D slope for the same reason that the 2D function ${\cal R}(M)$ is larger than the 3D function, namely the larger value of $N(M_{\rm min})^{1/2}$ than $N(M_{\rm min})^{1/3}$.
Figure 10: The relative average minimum separation is shown as a function of cluster mass, as in Figure 5, but now considering only cluster pairs at the same distance from the center of the cluster distribution, to remove the radial gradient of cluster mass shown in Fig. 9. The slope is still slightly negtive.
@@
The global distribution of clusters is highly irregular in this interacting pair with the center of the cluster coordinates moving nearly a kpc in distance while the interaction is taking place. Figure 8 shows the center at each Myr interval. The segregation of massive clusters illustrated by the negative slopes in Figures 5 and 6 corresponds to a centralization within the entire distribution of clusters, which, as seen in Figure 8, is distinct from a centralization in the galaxy-pair relative to the dark matter center of mass. Figure 9 shows the average cluster mass versus distance from the instantaneous center of the cluster distribution, i.e., from the points in Figure 8. There is a decreasing trend showing a segregation of more massive clusters toward this center. To check whether the entire segregation effect shown by ${\cal R}(M)$ is the result of this radial mass dependence, Figure 10 shows a modified ${\cal R}(M)$ function where only clusters considered within the same radial interval are used to search for near neighbors. That is, the radial range is divided into 500 pc bins (0-500pc, 500pc-1000pc, etc.) and near neighbors for each cluster are searched entirely within its bin, and also within a maximum distance proportional to the bin radius, namely within 250pc for the 0-500 pc bin, within 750 pc for the 500-1000pc bin, etc., to prevent cluster pairs from opposite sides of the galaxy. The result in Figure 10 still shows a slight negative slope, which implies that the mass segregation is somewhat three-dimensional.
@@ 5 Cluster Coalescence
Figure 12: The summed mass of a cluster pair is shown versus time in the simulation for all cluster pairs that are nearest neighbors regardless of mass, and which also have a ratio of relative kinetic energy to potential energy less than 1. Such near-neighbor pairs are likely to coalesce. Most of the thin horizontal streaks are single cluster pairs that have $KE/PE<1$ for the entire time of the streak; the clusters coalesce at the right-hand end of the streak. Compared to all nearest-neighbor cluster pairs, those with $KE/PE<1$ represent a fraction of 8.3%. The cyan curve represents the total cluster mass with $KE/PE<1$ as a function of time, using the right-hand axis.
To track such incidences in more detail, we plot in Figure 12 the summed mass versus the time for all cluster pairs that have $KE/PE<1$, regardless of mass. Each point in the plot indicates a cluster pair that is likely to coalesce. The points produce thin horizontal streaks in this plot when cluster pairs maintain $KE/PE<1$ over a period of time. The last time in a streak corresponds either to coalescence of the clusters or disruption of the pair. The cyan curve in the figure is the summed mass of clusters with $KE/PE<1$ as a function of time, using the right-hand ordinate for the scale. The number of nearest-cluster pairs with $KE/PE<1$ in the figure is 8821, which is 8.3% of the total number of nearest-cluster pairs at any $KE/PE$ value in all timesteps. If $KE/PE<1$ is an indication of eventual coalescence, then about this fraction of clusters coalesce.
The fractional mass of cluster pairs with $KE/PE<1$ is shown versus the SFR and CFE in Figure 13. The fractional mass of all cluster pairs, regardless of cluster age, is shown in blue, and the fractional mass for just the youngest clusters, where both members of the pair have ages less than 10 Myr, is shown in red. The lines in the plots connect the values calculated with 10 Myr averages, and the dots are for 2 Myr averages, showing the temporal scatter around the blue lines. The mass fraction susceptible to coalescence ($KE/PE<1$) is essentially zero for SFRs less than $\sim 10^{4}\;M_{\odot}$ Myr-1 and then it rises rapidly with SFR, reaching a peak mass fraction of about 0.6 at $10^{5}\;M_{\odot}$ Myr-1. For the young clusters (red dots), the fractional mass tracks the SFR above this limit. Note that the mass fraction of $KE/PE<1$ cluster pairs is much larger than the number fraction of $KE/PE<1$ cluster pairs, which was discussed in the previous paragraph. This is because massive clusters are more likely to coalesce than low-mass clusters (Fig. 11).
@@ 7 Conclusions
Star clusters that form in the dwarf galaxy collision simulated by Lahén et al. (2020) have the usual power-law mass function with slope $\alpha\sim-0.8$ to $-1$ from the time of their birth, with a maximum cluster mass that scales with the instantaneous SFR. The clusters also segregate in mass from the time of their birth, with more massive clusters centrally concentrated relative to each other than less massive clusters. Near-neighbor clusters are generally moving too fast to deflect by gravitational forces, but $\sim 8$% of them become significantly attracted to each other during the simulation, with a relative kinetic energy less than their potential energy. These 8% could coalesce. They represent up to 60% of the cluster mass at the peak of the SFR.
The implications of these results are that the mass distribution function and the spatial segregation of star clusters according to their masses are products of their birth that resist disruption for at least several hundred Myr. These properties are continuously present as new clusters appear, suggesting that the conditions for cluster formation, such as the intrinsic structure and motion of interstellar gas, are always present, regardless of how chaotic these conditions get from the galaxy collision and stellar feedback. Hierarchical cluster assembly has further implications for the formation of supermassive stars in cluster cores, while cluster mass segregation could promote the formation of nuclear star clusters.

### v2

@@ Abstract
The collective properties of star clusters are investigated using a simulation of the collision between two dwarf galaxies. The characteristic power law of the cluster mass function, $N(M)$, with a logarithmic slope $d\log N/d\log M\sim-1$, is present from cluster birth and remains throughout the simulation. The maximum mass of a young cluster scales with the star formation rate (SFR). The relative average minimum separation, ${\cal R}(M)=N(M)^{1/p}{\bar{D}}_{\rm min}(M)/{\bar{D}}(M_{\rm low})$, for average minimum distance ${\bar{D}}_{\rm min}(M)$ between clusters of mass $M$, and for lowest mass, $M_{\rm low}$, measured in projection ($p=2$) or three dimensions ($p=3$), has a negative slope, $d\log{\cal R}/d\log M\sim-0.2$, for all masses and ages. This agrees with observations of ${\cal R}(M)$ in low-mass galaxies studied previously. Like the slope of $N(M)$, ${\cal R}(M)$ is apparently a property of cluster birth for dwarf galaxies that does not depend on SFR or time. The negative slope for ${\cal R}(M)$ implies that massive clusters are more concentrated relative to lower mass clusters throughout the entire mass range. Cluster growth through coalescence is also investigated. The ratio of the kinetic to potential energy of all near-neighbor clusters is generally large, but a tail of low values in the distribution of this ratio suggests that a fraction of the clusters merge, $\sim 8$% by number throughout the $\sim 300$ Myr of the simulation and up to 60% by mass for young clusters in their first 10 Myr, scaling with the SFR above a certain threshold.
@@ 1 Introduction
Stars form in dense interstellar gas that has a hierarchical structure from turbulence and self-gravity (e.g., Scalo, 1985; Stanimirovic et al., 1999; Elmegreen et al., 2001). Young stellar positions are similarly hierarchical with star clusters forming in the mixed inner parts of the densest regions (e.g., Zhang et al., 2001; Elmegreen, 2010; Guszejnov et al., 2022). For a self-similar hierarchy, the mass distribution function of concentrations has a power law with a slope of around $-2$ ($-1$ on a log-log plot; Fleck, 1996; Elmegreen & Falgarone, 1996), which explains the power law portion of the mass function that is commonly observed for clusters (Elmegreen & Efremov, 1997; Zhang et al., 1999; Krumholz et al., 2019; Cook et al., 2019). The power-law portion of the stellar mass function may have a similar origin (Elmegreen, 1997; Hennebelle & Chabrier, 2008; Shadmehri & Elmegreen, 2011).
Power laws from structure rather than self-regulated processes have the property that the maximum mass of an object that forms scales with the total number of objects (Elmegreen, 1983). Such scaling has been observed for both star clusters (Larsen, 2002; Billett et al., 2002; Whitmore, 2002; Hunter et al., 2003; Bastian, 2008; Cook et al., 2023) and stars (Corbelli et al., 2009; Andrews et al., 2014; Jung et al., 2023). Physical models for the dependence of maximum cluster mass on star formation rate are in Weidner et al. (2004) and Randriamanakoto et al. (2013).
@@ 3 Cluster Mass Functions
Figure 2: The mass functions of clusters at the 5 fiducial times in increasing order of the colors: blue, red, brown, black, gray. The mass functions for all clusters are on the left and for clusters younger than 10 Myr are on the right. Power-law fits to the declining parts of the functions are shown in the same colors. The logarithmic slopes are all around $-1$, as indicated in Table 1.
@@
Figure 2 plots the cluster mass functions at the five fiducial times in the following sequence of colors: blue for the first time, 78 Myr, followed by red (128 Myr), brown (169 Myr), black (211 Myr), and gray for the last time, 277 Myr. The mass functions for all the clusters regardless of age are on the left and for clusters younger than 10 Myr are on the right. The straight lines are linear fits (on these log-log plots) to the right of the peaks. The slopes $\alpha$ of these lines are the power law indices of the mass functions. They are tabulated in Table 1. For a mass function written in linear intervals of mass, the slope is 1 less than $\alpha$, i.e., $n(M)\propto M^{\alpha-1}$ for masses between $M$ and $M+dM$ (i.e., $dn(M)/dM\propto M^{\alpha-1}$ and $d\log N(M)/d\log M\propto M^{\alpha}$). The highest-mass cluster is the single peak on the far right of these plots; it appears around the time of the peak in the SFR.
@@
Time (Myr) Slope [^2] These are slopes $\alpha$ on a $\log-\log$ plot where $-1$ is the conventional result. Subtract 1 to get the slope of the power law for the mass function itself, $n(M)\propto M^{\alpha-1}$. intercept
@@
Figure 3: The logarithmic slopes $\alpha$ of the mass functions for all clusters (blue) and young clusters (red), are shown versus time. They hover around $\alpha=-0.8$ to $-1$. The mass function itself, $n(M)dM$, has a slope that is more negative by 1.
@@
Figure 6: The logarithmic slope of the power-law relation between the relative average minimum separation and the cluster mass, $d\log{\cal R}(M)/d\log M$, is shown versus time for all clusters (top) and young clusters (bottom) and for projected, 2D (red) and 3D (blue) separations. The slope hovers around $-0.2$.
For a logarithmic mass function slope of $\alpha=-1$, the largest cluster mass is proportional to the number of clusters. Figure 4 shows the maximum cluster mass versus time in the left-hand panel, total stellar and cluster mass in the middle panel, and SFR in the right-hand panel. In the left and right-hand panels, the blue curves are for all cluster ages and the red curves are for clusters younger than 10 Myr. The maximum mass of a young cluster follows the SFR even as the SFR increases and decreases. The maximum mass for all ages does not follow the varying SFR because this mass does not decrease when the SFR drops.
@@ 4 Relative Average Minimum Separation and Projected Separation
Figure 7: The logarithmic slope of the power-law relation between the relative average minimum separation and the cluster mass, $d\log{\cal R}(M)/d\log M$, is shown versus cluster age for the time interval between 250 Myr and 299 Myr. Two-dimensional and three-dimensional cluster separations are shown in red and blue.
@@
Figure 5 shows ${\cal R}(M)$ for 3D (top) and 2D (bottom) distributions, and for all clusters (left) and clusters younger than 10 Myr (right). In each panel, ${\cal R}(M)$ is given at the 5 fiducial times with the same color code as in Figure 2. ${\cal R}(M)$ decreases with mass once there are enough clusters to obtain a meaningful slope, which is starting with the main starburst and for times of 169 Myr or later in the figure. Because both 3D and 2D curves have their minimum average separations ${\bar{D}}$ normalized at the lowest masses, their absolute values in these plots are proportional to $N(M)$ to the corresponding powers $1/3$ or $1/2$, making the 2D curves higher. The absolute value also depends on the mass interval used for the cluster counts, which is logarithmic in the figure, $d\log M=0.2$.
@@
Figure 6 shows the logarithmic slope $d\log{\cal R}(M)/d\log M$ versus time for 3D (blue) and 2D (red) distributions and for all clusters (top) and young clusters (bottom). Once there are enough clusters to have a reasonable slope, this slope is clearly negative even for young clusters. The 2D slope is slightly steeper than the 3D slope for the same reason that the 2D function ${\cal R}(M)$ is larger than the 3D function, namely the larger value of $N(M_{\rm min})^{1/2}$ than $N(M_{\rm min})^{1/3}$.
Figure 10: The relative average minimum separation is shown as a function of cluster mass, as in Figure 5, but now considering only cluster pairs at the same distance from the center of the cluster distribution, to remove the radial gradient of cluster mass shown in Fig. 9. The slope is still slightly negative.
@@
The global distribution of clusters is highly irregular in this interacting pair with the center of the cluster coordinates moving nearly a kpc in distance while the interaction is taking place. Figure 8 shows the center at each Myr interval. This irregularity is the result of the galaxy merger and is not a general feature of local dwarf galaxies. It is useful for the present study because we can view the cluster distribution independently of the galactic radial distribution, which cannot be done for symmetric galaxies. Nevertheless, there is an example of massive clusters together in a compact region, perhaps showing mass segregation relative to other clusters, in a lensed dwarf at redshift $\sim 10.2$ (Adamo et al., 2024).
The segregation of massive clusters illustrated by the negative slopes in Figures 5 and 6 corresponds to a centralization within the entire distribution of clusters, which, as seen in Figure 8, is distinct from a centralization in the galaxy-pair relative to the dark matter center of mass. Figure 9 shows the average cluster mass versus distance from the instantaneous center of the cluster distribution, i.e., from the points in Figure 8. There is a decreasing trend showing a segregation of more massive clusters toward this center. To check whether the entire segregation effect shown by ${\cal R}(M)$ is the result of this radial mass dependence, Figure 10 shows a modified ${\cal R}(M)$ function where only clusters considered within the same radial interval are used to search for near neighbors. That is, the radial range is divided into 500 pc bins (0-500pc, 500pc-1000pc, etc.) and near neighbors for each cluster are searched entirely within its bin, and also within a maximum distance proportional to the bin radius, namely within 250pc for the 0-500 pc bin, within 750 pc for the 500-1000pc bin, etc., to prevent cluster pairs from opposite sides of the galaxy. The result in Figure 10 shows the trend. The fits have logarithmic slopes of $-0.15\pm 1.09$, $-0.10\pm 0.16$, $-0.27\pm 0.25$, $-0.09\pm 0.15$ for the four times, in order.
@@ 5 Cluster Coalescence
Figure 12: The summed mass of a cluster pair is shown versus time in the simulation for all cluster pairs that are nearest neighbors regardless of mass, and which also have a ratio of relative kinetic energy to potential energy less than 1. Such near-neighbor pairs are likely to coalesce. Most of the thin horizontal streaks are single cluster pairs that have $KE/PE<1$ for the entire time of the streak; the clusters coalesce at the right-hand end of the streak. Compared to all nearest-neighbor cluster pairs, those with $KE/PE<1$ represent a fraction of 8.3%. The red curve represents the total cluster mass with $KE/PE<1$ as a function of time, using the right-hand axis.
To track such incidences in more detail, we plot in Figure 12 the summed mass of clusters in a pair versus the time for all cluster pairs that have $KE/PE<1$, regardless of mass. Each point in the plot indicates a cluster pair that is likely to coalesce. The points produce thin horizontal streaks in this plot when cluster pairs maintain $KE/PE<1$ over a period of time. The last time in a streak corresponds either to coalescence of the clusters or disruption of the pair. The red curve in the figure is the summed mass of all clusters with $KE/PE<1$ as a function of time, using the right-hand ordinate for the scale. The number of nearest-cluster pairs with $KE/PE<1$ in the figure is 8821, which is 8.3% of the total number of nearest-cluster pairs at any $KE/PE$ value in all timesteps. If $KE/PE<1$ is an indication of eventual coalescence, then about this fraction of clusters coalesce.
The fractional mass of cluster pairs with $KE/PE<1$ is shown versus the SFR and CFE in Figure 13. The fractional mass of all cluster pairs, regardless of cluster age, is shown in blue, and the fractional mass for just the youngest clusters, where both members of the pair have ages less than 10 Myr, is shown in red. The lines in the plots connect the values calculated with 10 Myr averages, and the dots are for 2 Myr averages, showing the temporal scatter around the lines. The mass fraction susceptible to coalescence ($KE/PE<1$) is essentially zero for SFRs less than $\sim 10^{4}\;M_{\odot}$ Myr-1 and then it rises rapidly with SFR, reaching a peak mass fraction of about 0.6 at $10^{5}\;M_{\odot}$ Myr-1. For the young clusters (red dots), the fractional mass tracks the SFR above this limit. Note that the mass fraction of $KE/PE<1$ cluster pairs is much larger than the number fraction of $KE/PE<1$ cluster pairs, which was discussed in the previous paragraph. This is because massive clusters are more likely to coalesce than low-mass clusters (Fig. 11).
@@ 7 Conclusions
Star clusters that form in the dwarf galaxy collision simulated by Lahén et al. (2020) have the usual power-law mass function with logarithmic slope $\alpha\sim-0.8$ to $-1$ from the time of their birth, with a maximum cluster mass that scales with the instantaneous SFR. The clusters also segregate in mass from the time of their birth, with more massive clusters centrally concentrated relative to each other than less massive clusters. Near-neighbor clusters are generally moving too fast to deflect by gravitational forces, but $\sim 8$% of them become significantly attracted to each other during the simulation, with a relative kinetic energy less than their potential energy. These 8% could coalesce. They represent up to 60% of the cluster mass at the peak of the SFR.
The implications of these results are that the mass distribution function and the spatial segregation of star clusters according to their masses are products of their birth that resist disruption for at least several hundred Myr. These properties are continuously present as new clusters appear, suggesting that the conditions for cluster formation, such as the intrinsic structure and motion of interstellar gas, are always present, regardless of how chaotic these conditions get from the galaxy collision and stellar feedback. The preferred centralized location of massive clusters demonstrates either an environmental effect for cluster mass or a migration of massive clusters soon after they form. Environmental effects add physical processes to the determination of cluster mass, in addition to purely stochastic sampling of the mass function (Pflamm-Altenburg et al., 2013). Hierarchical cluster assembly has further implications for the formation of supermassive stars in cluster cores, while cluster mass segregation could promote the formation of nuclear star clusters.